# Pluto and Charon: A Time and Memory Efficient Collaborative Edge AI Framework for Personal LLMs Fine-Tuning

### v1

@@ 5.1. Resource-Efficient Collaborative Orchestration for LLMs Fine-Tuning
Upon the completion of dynamic programming process, we obtain a set of balanced partition configurations for various number of pipeline stages: $\{W_{s}|\text{ config. of }W(0\rightarrow L,\mathcal{D},s),s\in\{1,2,...,|\mathcal{D}|\}\}$. The next step is to determine the optimal number of stages. Using recorded configurations, we can profile FP and BP execution time of stage $i$ in $W_{s}$ as $e_{f}^{s}(i)$ and $e_{b}^{s}(i)$. Similarly, forward and backward communication time between stages $i$ and $i+1$ are represented as $c_{f}^{s}(i)$ and $c_{b}^{s}(i)$. $\text{AR}^{s}(i)$ represents the AllReduce time of stage $i$ in $W_{s}$. As shown in Figure 6(b), we can divide per mini-batch training of $W_{s}$ into three phases: beginning phase, execution phase, and ending phase with corresponding times denoted as $L_{b}^{s},L_{e}^{s},L_{n}^{s}$:
(4) $\vspace{-15pt}L_{b}^{s}=\sum_{i=1}^{s-1}[e_{f}^{s}(i)+c_{f}^{s}(i)],\quad L_{e}^{s}=M\cdot(e_{f}^{s}(s)+e_{b}^{s}(s)),$
(5) $L_{n}^{s}=\max_{i\in\{1...,s\}}(\text{AR}^{s}(i)+\displaystyle\sum_{j=i}^{s-1}(e_{b}^{s}(j)+c_{b}^{s}(j)),$
(6) $\min_{s}{(L_{b}^{s}+L_{e}^{s}+L_{n}^{s})}.\vspace{-15pt}$
@@
Our algorithm aims to minimize this total latency by optimally determining the number of stages $s$. We remark that our parallelism planning is an offline procedure that runs once before deployment. The time complexity for our dynamic programming algorithm exhibits an upper bound of $O(L^{2}|\mathcal{D}|^{2})$. In our experiment, the whole planning time is within three seconds on an edge device.
@@
5.2. Cache-Enabled Collaborative Edge Fine-Tuning of Parallel Adapters
@@
Data-Parallel Fine-Tuning for Parallel Adapters The computationally lightweight nature of the Parallel Adapters precludes the use of pipeline parallelism to fine-tuning with activation cache, as it would result in unoverlapable inter-stage communication latency. Therefore, we employ data parallelism to exclusively fine-tune our Parallel Adapters. Specifically, after the first training epoch, the activation cache for all samples is already collected. We then perform collective communication to redistribute the Parallel Adapters parameters and locally cached activations across all devices, ensuring each device receives the complete set of adapter parameters and corresponding activations. The devices then utilize this shared information to fine-tune the parallel adapters in a data-parallel manner. In our experiments, fine-tuning the BART-Large model on the MRPC dataset for three epochs, the redistribution of parameters and activations only contributed to approximately $8\%$ of the total training time. Notably, the overhead of this process can be further amortized over additional training epochs. An instance of personal LLMs fine-tuning with activation cache is depicted in Figure 7.
Figure 7. An instance of fine-tuning with activation cache.
@@
Table 2. Training durations (in hours) for different methods: 3 epochs for MRPC and STS-B, and 1 epoch for SST-2 and QNLI.
@@
Eco-FL 0.41 0.64 2.45 3.87 0.55 0.87 3.33 5.26 2.73 4.28 16.48 26.02
EDDL 0.31 0.48 1.86 2.94 OOM OOM OOM OOM OOM OOM OOM OOM
Parallel Adapters PAC (Ours) 0.14 0.22 1.34 2.12 0.29 0.45 2.69 4.25 0.69 1.09 8.88 14.02
Table 3. Comparison of final performance between different fine-tuning techniques across four datasets. We report the average of F1 score and accuracy for MRPC. We use Pearson-Spearman Correlation as the metric for STS-B. For SST-2 and QNLI, we report accuracy. The mean value is the average performance of Full Model, Adapters and LoRA.
@@
Parallel Adapters (Ours) 88.24 90.43 93.46 93.25 87.71 90.54 95.25 93.68 91.7 91.57 95.76 93.7
@@
6. Evaluation
6.1. Implementation and Setups
Implementation of PAC. We have fully implemented the prototype framework of PAC and baselines with $\sim$2,00020002,0002 , 000 LoC in Python atop Pytorch (pyt, 2019). PAC’s idea is also portable and can work well with other lightweight ML frameworks such as MNN (Jiang et al., 2020) and TF-Lite (tfl, 2021). Our Parallel Adapters is a lightweight version of the backbone model. The size of Parallel Adapters is determined by the reduction factor $k$. All weights and hidden state dimensions of the Parallel Adapters are $\frac{1}{k}$ times the corresponding weights and hidden states of the backbone model. In our experiments, the reduction factor $k$ is set to 8. The weights of the Parallel Adapters are initialized based on structural pruning, using the weights of the backbone model. We insert Parallel Adapters at the end of each transformer layer.
Models and Datasets. We evaluate PAC with three typical transformer based LLM with parameters ranging from $0.25$B to $0.74$B, as detailed in Table 4, which are widely considered for IPA and edge deployments (Li et al., 2024; Yuan et al., 2023). All experiments were performed under conditions using Float32 precision to ensure fine-tuning performance. We employ two variants of the T5 model (Raffel et al., 2020), specifically T5-Base and T5-Large with differing parameter sizes. We also compare PAC with baseline methods with BART-Large (Lewis et al., 2019) as the backbone for our parallel adapters. We evaluate our fine-tuned LLMs with four tasks from GLUE benchmark. The four tasks evaluate models on multiple diverse tasks over sentiment analysis (SST2), similarity and paraphrase (MRPC, STS-B) and natural language inference (QNLI).
Edge Environment Setup. We evaluate PAC across a realistic edge platform consisting of multiple NVIDIA Jetson Nano (jet, 2019), widely recognized as prevalent off-the-shelf edge devices. Each device is equipped with a $128$-core NVIDIA Maxwell GPU and 4GB unified memory. We simulate common network conditions in edge environments (e.g., smart homes) by setting the intra-cluster network bandwidth to $1000$Mbps.
Baseline Methods. We compare PAC with both single-device method and the state-of-the-art collaborative edge training methods: (1) Standalone means fine-tuning LLMs on a single edge device. We compare with it to analyze the scalability performance of PAC. (2) Eco-FL (Ye et al., 2022) is a collaborative edge system that facilitates pipeline parallelism training across an edge device cluster within the same local area network, segmenting LLMs into sequential stages for processing in a pipeline fashion. (3) EDDL (Hao and Zhang, 2021) employs conventional data parallel training across edge devices, distributing batch data among cluster devices for simultaneous processing.
Table 4. LLM model specifications used for experiments. "en-de" indicates encoder-decoder LLM structure.
Model Structure Layers Heads
Hidden
Size
Param.
Count
T5-Base (Raffel et al., 2020)
en-de
12 12 768 0.25B
BART-Large (Lewis et al., 2019)
en-de
12 16 1024 0.41B
T5-Large (Raffel et al., 2020)
en-de
24 16 1024 0.74B
Considering that the aforementioned baseline systems were not specifically designed for the fine-tuning of LLMs, we ensure a fair comparison by equipping these edge systems with various LLM fine-tuning techniques. These include full model fine-tuning and popular PEFT techniques. (1) In Full model fine-tuning, all the LLM parameters are updated for a downstream task. (2) LoRA (Hu et al., 2021) is a widely-used PEFT technique that decomposes the parameter update for a weight matrix into two trainable low-rank matrices. (3) Adapters (Houlsby et al., 2019) is another widely-used PEFT technique that injects small trainable modules at the end of each transformer layer.
6.2. End-to-end Performance
Table 2 and Table 3 summarize the end-to-end performance comparisons between PAC, the single-device method, and state-of-the-art collaborative edge training methods. To ensure fair comparisons, these baseline methods are enhanced with prevalent PEFT techniques, including Adapters and LoRA. Fine-tuning the smaller datasets, MRPC and STS-B, is conducted over three epochs, with the latter two epochs benefiting from the PAC activation cache. Conversely, for larger datasets such as STS-2 and QNLI, a single epoch of fine-tuning is sufficient to achieve satisfactory performance.
@@ 6.2. End-to-end Performance
Figure 8. Comparison of different fine-tuning techniques. P.A. indicates our Parallel Adapters technique. Mini-batch size 16; sequence length: 128.
PAC significantly speeds up the training process while preserving convergence performance. PAC achieves an acceleration ranging from 1.21×1.21$\times$1.21 × to 5.44×5.$4$ on SST-2 and QNLI without utilizing activation cache. In comparison to Standalone and EDDL, these two baselines often encounter Out-of-Memory (OOM) issues, even when integrating PEFT techniques such as LoRA and Adapters. This issue stems from the training requirement for each edge device to host the entire target model. Particularly for T5-Large, a single Jetson Nano is inadequate to accommodate LLM parameters, not to mention the intermediate activations. Compared to Eco-FL, PAC achieves an acceleration ranging from 1.21×1.21$\times$1.21 × to 5.41×5.41$\times$5.41 × on SST-2 and QNLI, without utilizing activation cache. Parallel Adapters not only alleviate the memory footprint of LLM parameters but also intermediate activations. Eco-FL’s pipeline parallel strategies allow each edge device to host only a portion of the model parameters. However, these devices still bear a substantial memory footprint from intermediate activations, even when employing PEFT technologies such as LoRA and Adapters. Therefore, the Eco-FL approach necessitates the use of smaller micro-batch sizes or a reduction in the number of micro-batches simultaneously input into the pipeline. This results in decreased concurrency in pipeline parallelism and lowers the training throughput. Moreover, our hybrid parallelism merges the benefits of both data and pipeline parallelism, providing an expanded search space for parallel architectures to accommodate complex edge environments. Our method enables the identification of the most efficient parallel configuration with maximum throughput within the constraints of available resources. With the integration of our activation cache mechanism, PAC achieves speedups of up to 8.64×8.64$\times$8.64 × on the MRPC and STS-B datasets. As discussed in §5, our Parallel Adapters constitute a lightweight, independent network. We can skip both the forward and backward passes through the LLM backbone, since the required activations have already been calculated and stored. Consequently, training overhead can be markedly reduced in the second and third fine-tuning epochs.
Table 3 displays the performance of various full model and PEFT fine-tuning methods on four datasets after training. Fine-tuning involves 3 epochs for the smaller MRPC and STS-B datasets, and 1 epoch for the larger SST-2 and QNLI datasets. We can observe that PAC achieves comparable or superior performance to full model fine-tuning and PEFT techniques across various models and datasets. The largest discrepancy in mean performance metrics between PAC and these methods is only -0.37, a negligible difference. Notably, PAC frequently outperforms these methods and achieves the highest performance on the SST-2 dataset with the T5-Large model.

### v2

@@ V-A Resource-Efficient Collaborative Orchestration for LLMs Fine-Tuning
Upon the completion of dynamic programming process, we obtain a set of balanced partition configurations for various number of pipeline stages: $\{W_{s}|\text{ config. of }W(0\rightarrow L,\mathcal{D},s),s\in\{1,2,...,|\mathcal{D}|\}\}$. The next step is to determine the optimal number of stages. Using recorded configurations, we can profile FP and BP execution time of stage $i$ in $W_{s}$ as $e_{f}^{s}(i)$ and $e_{b}^{s}(i)$. Similarly, forward and backward communication time between stages $i$ and $i+1$ are represented as $c_{f}^{s}(i)$ and $c_{b}^{s}(i)$. $\text{AR}^{s}(i)$ represents the AllReduce time of stage $i$ in $W_{s}$. $M$ is the number of micro-batch. As shown in Figure 10(b), we can divide per mini-batch training of $W_{s}$ into three phases: beginning phase, execution phase, and ending phase with corresponding latency denoted as $L_{b}^{s},L_{e}^{s},L_{n}^{s}$:
$\vskip-15.0ptL_{b}^{s}=\sum_{i=1}^{s-1}[e_{f}^{s}(i)+c_{f}^{s}(i)],\quad L_{e}^{s}=M\cdot(e_{f}^{s}(s)+e_{b}^{s}(s)),$ (5)
$L_{n}^{s}=\max_{i\in\{1...,s\}}(\text{AR}^{s}(i)+\displaystyle\sum_{j=i}^{s-1}(e_{b}^{s}(j)+c_{b}^{s}(j)),$ (6)
$\sigma=\operatorname*{argmin}_{s}(L_{b}^{s}+L_{e}^{s}+L_{n}^{s}).$ (7)
@@
Our algorithm aims to minimize this total latency by optimally determining the number of stages $s$. We summarize our dynamic programming planning in Algorithm 1. We remark that our parallelism planning is an offline procedure that runs once before deployment. The time complexity for our dynamic programming algorithm exhibits an upper bound of $O(BL^{2}|\mathcal{D}|^{3})$. In our experiment, the whole planning time is within several minutes on an edge device.
@@
V-B Cache-Enabled Collaborative Edge Fine-Tuning of Parallel Adapters
@@
Data-Parallel Fine-Tuning for Parallel Adapters. The computationally lightweight nature of the Parallel Adapters precludes the use of pipeline parallelism to fine-tuning with activation cache, as it would result in unoverlapable inter-stage communication latency. Therefore, we employ data parallelism to exclusively fine-tune our Parallel Adapters. Specifically, after the first training epoch, the activation cache for all samples is already collected. We then perform collective communication to redistribute the Parallel Adapters parameters and locally cached activations across all devices, ensuring each device receives the complete set of adapter parameters and corresponding activations. The devices then utilize this shared information to fine-tune the parallel adapters in a data-parallel manner. In our experiments, fine-tuning the BART-Large model on the MRPC dataset for three epochs, the redistribution of parameters and activations only contributed to approximately $8\%$ of the total training time. Notably, the overhead of this process can be further amortized over additional training epochs. An instance of personal LLMs fine-tuning with activation cache is depicted in Figure 11.
Figure 11: An instance of fine-tuning with activation cache.
@@
VI Evaluation
VI-A Implementation and Setups
Implementation of PAC+. We have fully implemented the prototype framework of PAC+ and baselines with $\sim$$2,000$ LoC in Python atop Pytorch [35]. PAC+’s idea is also portable and can work well with other lightweight ML frameworks such as MNN [20] and TF-Lite [31]. Our Parallel Adapters is a lightweight version of the backbone model. The size of Parallel Adapters is determined by the reduction factor $r$. All weights and hidden state dimensions of the Parallel Adapters are $\frac{1}{r}$ times the corresponding weights and hidden states of the backbone model. In our experiments, the reduction factor $r$ is set to 8. The weights of the Parallel Adapters are initialized based on structural pruning, using the weights of the backbone model. We insert Parallel Adapters at the end of each transformer layer.
Models and Datasets. We evaluate PAC+ with three typical transformer based LLM with parameters ranging from 0.250.25B to 0.740.74B, as detailed in Table III, which are widely considered for IPA and edge deployments [25, 49]. All experiments were performed under conditions using FP32 precision to ensure fine-tuning performance. We employ two variants of the T5 model [36], specifically T5-Base and T5-Large with differing parameter sizes. We also compare PAC+ with baseline methods with BART-Large [22] as the backbone for our parallel adapters. We evaluate our fine-tuned LLMs with four tasks from GLUE benchmark. The four tasks evaluate models on multiple diverse tasks over sentiment analysis (SST2), similarity and paraphrase (MRPC, STS-B) and natural language inference (QNLI).
TABLE III: LLM model specifications used for experiments. "en-de" indicates encoder-decoder LLM structure.
Model Structure Layers Heads
Hidden
Size
Param.
Count
T5-Base [36]
en-de
12 12 768 0.25B
BART-Large [22]
en-de
12 16 1024 0.41B
T5-Large [36]
en-de
24 16 1024 0.74B
TABLE IV: Specifications of edge devices in Evaluation.
Edge Device GPU Processor
Memory
Budget
GPU
Frequency
Power
Mode
Jetson Nano [18] 128-core Maxwell 4GB 921MHz 10W (H)
640MHz 5W (L)
Jetson TX2 [19] 256-core Pascal 8GB 1.3GHz 15W (H)
850MHz 7.5W (L)
TABLE V: Training durations (in hours) for different methods: 3 epochs for MRPC and STS-B, and 1 epoch for SST-2 and QNLI.
@@ VI-A Implementation and Setups
PP (Eco-FL) 0.41 0.64 2.45 3.87 0.55 0.87 3.33 5.26 2.73 4.28 16.48 26.02
DP (EDDL) 0.31 0.48 1.86 2.94 OOM OOM OOM OOM OOM OOM OOM OOM
Parallel Adapters PAC+ (Ours) 0.14 0.22 1.34 2.12 0.29 0.45 2.69 4.25 0.69 1.09 8.88 14.02
TABLE VI: Comparison of final performance between different fine-tuning techniques across four datasets. We report the average of F1 score and accuracy for MRPC. We use Pearson-Spearman Correlation as the metric for STS-B. For SST-2 and QNLI, we report accuracy. The mean value is the average performance of Full Model, Adapters and LoRA.
@@
Parallel Adapters (Ours) 88.24 90.43 93.46 93.25 87.71 90.54 95.25 93.68 91.70 91.57 95.76 93.70
@@
Edge Environment Setup. We conduct experiments using two off-the-shelf edge devices listed in Table IV, simulating four heterogeneous platforms by adjusting their power modes. We evaluate PAC+’s performance in two realistic edge environments, incorporating both homogeneous and heterogeneous configurations. Homogeneous Environment A (Env. A) consists of $4\timesNano$-H, while Heterogeneous Environment B (Env. B) comprises $1\timesNano$-H, $1\timesNano$-L, $1\timesTX$-H and $1\timesTX$-L. We simulate common network conditions in edge environments (e.g., smart homes) by setting the intra-cluster network bandwidth to 10001000Mbps.
Baseline Methods. To evaluate the high performance and robustness of our PAC+ system, we comprehensively compare it with a wide range of baseline methods, including those we build ourselves and existing end-to-end collaborative edge training systems.
To thoroughly evaluate the advantages of our PAC+ in both algorithmic and system design, we select 3 classic collaborative edge computing paradigms and 3 widely adopted LLM fine-tuning methods, combining them in pairs to form 9 distinct baselines. The three collaborative edge computing paradigms are as follows:
  1. 1.
    Standalone means fine-tuning LLMs on a single edge device. We compare with it to analyze the scalability performance of PAC+.
  2. 2.
    Data Parallelism (DP) [24] is a classic cluster computing paradigm that distributes batch data across devices for simultaneous processing. EDDL [11] is a representative work that applies DP to edge computing, and we reference its architecture to implement our baseline.
  3. 3.
    Pipeline Parallelism ($P$) [17] enables collaborative training across an edge device cluster by segmenting LLMs into sequential stages for pipeline processing. We implement our baseline framework based on the representative work Eco-FL [47].
The three LLM fine-tuning methods are as follows:
  1. 1.
    Full model fine-tuning means that all the LLM parameters are updated for a downstream task.
  2. 2.
    LoRA [16] is a widely-used PEFT technique that decomposes the parameter update for a weight matrix into two trainable low-rank matrices. We apply LoRA to $W_{q}$ and $W_{v}$ following the settings from [16].
  3. 3.
    Adapters [13] is another widely-used PEFT technique that injects small trainable modules at the end of each transformer layer following the settings from [13].
In addition to our self-built baselines, we also compare PAC+ with two existing end-to-end edge training systems optimized for resource heterogeneity in edge environments:
  1. 1.
    Asteroid [46] is a collaborative DNN training framework designed for heterogeneous edge device clusters. It employs a hybrid data and pipeline parallelism architecture, similar to PAC+, but is designed for full-parameter fine-tuning of both vision and language models.
  2. 2.
    HetPipe [33] is a distributed training framework designed for heterogeneous consumer-grade GPU clusters. It facilitates asynchronous parallel training by treating sub-groups of GPUs as virtual workers, employing intra-worker PP and inter-worker communication.
VI-B Analysis of Algorithmic and System Design Benefits
(1) Comparison of End-to-End Elapsed Fine-Tuning Time Between PAC+ and Baselines. To thoroughly evaluate the strengths of PAC+ in both algorithmic and system design, we select three classic collaborative edge computing paradigms and equip them with three prevalent LLM fine-tuning techniques as baselines in the edge environment Env. A. For the Eco-FL, we divided one mini-batch into 4 micro-batches to enhance pipeline concurrency, whereas the Standalone and EDDL methods were fine-tuned strictly at the mini-batch granularity. Table V summarizes the end-to-end fine-tuning elapsed time on the GLUE datasets for both PAC+ and the baseline methods. For fine-tuning smaller datasets in GLUE, MRPC and STS-B, we conduct training over three epochs, with the latter two epochs benefiting from the PAC+ activation cache. In contrast, for larger datasets in GLUE, STS-2 and QNLI, a single epoch of fine-tuning is sufficient to achieve satisfactory performance.
As shown in Table V, PAC+ significantly speeds up the training process while preserving convergence performance. PAC+ achieves an acceleration ranging from $1.21\times$ to $5.44\times$ on SST-2 and QNLI without utilizing activation cache. In comparison to Standalone and DP, these two baselines often encounter Out-of-Memory (OOM) issues, even when integrating PEFT techniques such as LoRA and Adapters. This issue stems from the training requirement for each edge device to host the entire target model. Particularly for T5-Large, a single Jetson Nano is inadequate to accommodate LLM parameters, not to mention the intermediate activations. Compared to PP, PAC+ achieves an acceleration ranging from $1.21\times$ to $5.41\times$ on SST-2 and QNLI, without utilizing activation cache. Parallel Adapters not only alleviate the memory footprint of LLM parameters but also intermediate activations. Pipeline parallel strategies allow each edge device to host only a portion of the model parameters. However, these devices still bear a substantial memory footprint from intermediate activations, even when employing PEFT technologies such as LoRA and Adapters. Therefore, the PP approach necessitates the use of smaller micro-batch sizes or a reduction in the number of micro-batches simultaneously input into the pipeline. This results in decreased concurrency in pipeline parallelism and lowers the training throughput. Moreover, our hybrid parallelism merges the benefits of both data and pipeline parallelism, providing an expanded search space for parallel architectures to accommodate complex edge environments. Our method enables the identification of the most efficient parallel configuration with maximum throughput within the constraints of available resources. With the integration of our activation cache mechanism, PAC+ achieves speedups of up to $8.64\times$ on the MRPC and STS-B datasets. As discussed in §5, our Parallel Adapters constitute a lightweight, independent network. We can skip both the forward and backward passes through the LLM backbone, since the required activations have already been calculated and stored. Consequently, training overhead can be markedly reduced in the second and third fine-tuning epochs.
(2) Analysis of Fine-Tuned LLM Performance with PAC+. Table V demonstrates that our PAC+ achieves higher fine-tuning throughput compared to all baselines. However, it is still necessary to further validate that our PAC+ design does not compromise model accuracy in exchange for lower elapsed time. Table VI reports the model performance of LLMs fine-tuned using three classic LLM fine-tuning algorithms and our Parallel Adapters approach. All accuracy metrics were evaluated on a server environment with sufficient computational and memory resources to ensure consistent performance reporting, irrespective of potential OOM constraints. Fine-tuning involves 3 epochs for the smaller MRPC and STS-B datasets, and 1 epoch for the larger SST-2 and QNLI datasets. We can observe from Table VI that PAC+ achieves comparable or superior performance to full model fine-tuning and baseline PEFT techniques across various models and datasets. The largest discrepancy in mean performance metrics between PAC+ and these baseline methods is only -0.37, a negligible difference. Notably, PAC+ frequently outperforms these methods and achieves the highest performance on the SST-2 dataset with the T5-Large model.
(a) Total time required to fine-tune 1 epoch on the MRPC dataset in Env. B.
(b) Total time required to fine-tune 3 epoch on the MRPC dataset in Env. B. PAC+ further benefits from the activation cache design.
Figure 12: Comparison of total fine-tuning time across different baselines. PAC+ (Homo) refers to a version of the PAC+ algorithm that does not consider device resource heterogeneity.
VI-C Comparison with Existing Collaborative Training Systems under Heterogeneous Edge Environments
To evaluate the superiority and robustness of our proposed PAC+ system compared to existing collaborative edge training systems, we select two state-of-the-art training frameworks designed for heterogeneous device clusters, Asteroid [46] and HetPipe [33], as baselines for comparison with PAC+. The experiments are conducted in the heterogeneous edge environment Env. B. In addition to comparing PAC+ with the two baselines, we also included its older PAC version [32], which does not account for device heterogeneity, as part of our ablation study. This comparison aims to demonstrate the effectiveness of our heterogeneity-aware algorithm design. We evaluated all three models on the MRPC dataset, conducting fine-tuning for one epoch (Figure 12(a)) and three epochs (Figure 12(b)), respectively.
@@ VI-C Comparison with Existing Collaborative Training Systems under Heterogeneous Edge Environments
Figure 13: Comparison of different fine-tuning techniques. P.A. indicates our Parallel Adapters technique. Mini-batch size 16; sequence length: 128.
(a) Fine-tuning performance of different initialization strategies on BART-Large.
(b) Fine-tuning performance of different initialization strategies on T5-Large.
Figure 14: The ablation study on different initialization strategies for Parallel Adapters using MRPC dataset.
TABLE VII: Comparison of final performance between different quantization precision across four datasets. We report the average of F1 score and accuracy for MRPC. We use Pearson-Spearman Correlation as the metric for STS-B. For SST-2 and QNLI, we report accuracy. The mean value is the average performance of Full Model, Adapters and LoRA.
Backbone LLM Precision T5-Base BART-Large T5-Large
MRPC STS-B SST-2 QNLI MRPC STS-B SST-2 QNLI MRPC STS-B SST-2 QNLI
FP32 88.24 90.43 93.46 93.25 87.71 90.54 95.25 93.68 91.70 91.57 95.76 93.70
FP16 88.42 89.33 93.35 92.58 87.76 90.34 94.83 92.62 91.24 91.17 95.43 92.88
INT8 87.45 89.12 92.78 91.97 87.66 90.03 94.64 92.10 91.03 90.17 94.73 92.23
INT4 87.24 88.25 92.66 91.10 87.52 89.87 94.49 91.27 89.69 90.08 94.72 91.71
As shown in Figure 12(a), under the one-epoch fine-tuning setting, the PAC+ architecture achieves a $3.2\times$$9.7\times$ improvement over HetPipe and a $2.9\times$$8.1\times$ improvement over Asteroid, attributed to the co-design of the PAC+ algorithm and system. Specifically, HetPipe adopts a hybrid parallelism architecture with inter-group DP and intra-group PP, combining asynchronous updates to address device heterogeneity. However, this parallelism architecture has been shown in existing studies [46] to require higher inter-device communication, making it unsuitable for bandwidth-constrained edge clusters. Asteroid adopts the same inter-group PP and intra-group DP architecture as PAC+, along with fine-grained resource partitioning and orchestration to address heterogeneity. However, both Asteroid and HetPipe lacks optimizations for LLM fine-tuning algorithms and relies solely on full-parameter fine-tuning, resulting in higher computational resource requirements and longer training times. Furthermore, the activation cache design in PAC+ enables even greater performance improvements during three-epoch fine-tuning compared to the baselines, as shown in Figure 12(b). Specifically, PAC+ achieves a $7.6\times$$14.7\times$ improvement over HetPipe and a $6.9\times$$12.3\times$ improvement over Asteroid. Our designed PAC+ algorithm can achieve up to a $35\%$ latency reduction compared to algorithms that do not account for device heterogeneity.
VI-D Further Breakdown of the Benefits of Parallel Adapters
We further conduct experiments to break down the training latency and memory footprint, assessing the time and memory efficiency of Parallel Adapters at the edge. We conduct experiments on an edge cluster consisting of 8 Jetson Nano-H. All baselines are implemented using the same hybrid parallel architecture and equipped with different fine-tuning algorithms, with 1F1B micro-batch scheduling disabled. "Activations" contain the intermediate results and optimizer states. Figure 13 illustrates that Parallel Adapters outperform other fine-tuning techniques regarding both time and memory efficiency.
(1) Parallel Adapters markedly reduce per-sample training time. Figure 13(a) presents the average sample training time across different fine-tuning techniques. Without activation cache, Parallel Adapters can reduce the average sample training time by $31.94\%$ to $56.24\%$ compared to baseline methods, primarily owing to a substantial decrease in backward propagation overhead. Both Adapters and LoRA incorporate trainable structures into the backbone model, thus necessitating backpropagation across the entire backbone model for gradient computation of these parameters. Consequently, regarding backward time, Adapters and LoRA can only achieve approximately a $49\%$ reduction compared to full fine-tuning. In contrast, backpropagation through the backbone model is unnecessary with Parallel Adapters, leading to a more substantial reduction in backward time, nearly $92\%$ compared to full fine-tuning. Moreover, Parallel Adapters With activation cache mechanism can further decrease the average sample training time up to $96.39\%$. These results demonstrate the substantial reduction in training time achieved by Parallel Adapters.
Figure 15: Memory footprint analysis of baselines and Parallel Adapters with various quantization precision
(2) Parallel Adapters yield a substantial reduction in memory usage. Figure 13(b) depicts the breakdown of the memory footprint for different fine-tuning techniques. We report the peak memory consumption per device across edge clusters. Without activation cache, Parallel Adapters can reduce memory usage by $25.27\%$ to $60.49\%$. Adapters and LoRA demonstrate parameter efficiency but do not exhibit significant memory efficiency. While these techniques notably decrease the memory footprint of gradients by reducing the number of trainable parameters, the memory usage associated with activations remains considerable. However, intermediate activations often become the primary memory bottleneck during training, especially with larger batch sizes. For PEFT techniques such as Adapters and LoRA, activation memory can be reduced by up to $28.15\%$ compared to full fine-tuning across the three models. In contrast, Parallel Adapters achieve a more significant reduction, reaching up to $58.87\%$. With activation cache, Parallel Adapters can decrease the peak memory footprint from $74.57\%$ to $88.16\%$ compared to baselines. This is because it’s sufficient to store only the lightweight Parallel Adapters, eliminating the need to host the entire LLM backbone in memory.
VI-E Analyzing the Impact of Weight Initialization on Parallel Adapters
To investigate the impact of weight initialization on the fine-tuning performance of Parallel Adapters, we compare four initialization methods: knowledge distillation, structural pruning, random Gaussian initialization, and zero initialization. The fine-tuning performance of the BART-Large and T5-Large models on the MRPC dataset is illustrated in Figure 14. We implemented our distillation method using Google Research’s open-source distillation toolkit [14] and applied structural pruning with Torch-Pruning, a popular open-source toolkit for structural pruning built on PyTorch [6, 38]. Random Gaussian initialization is widely used in popular PEFT techniques, such as LoRA [16]. We can observe from Figure 14 that using distillation or pruning-based initialization for Parallel Adapters allows the model to reach the target accuracy more quickly during fine-tuning compared to random Gaussian or zero initialization. Specifically, using distillation and pruning-based initialization, the BART-Large model achieves a combined score of 0.875 after approximately 410 and 457 iterations, respectively, while random Gaussian initialization requires 605 iterations. Reducing the required fine-tuning iterations can significantly lower the time and computational resources required for fine-tuning on resource-constrained edge clusters.